OpenStudy (anonymous):

Muzzles A and B are 700 m apart and are placed in the ground. Muzzle A is projected 30 degrees above the horizontal and to the direction where Muzzle B is located. Muzzle B is projected vertically upward. Muzzle A fires a canon shell at the speed of 100 m/s. If muzzle B fires a canon shell at the same instant, what should be the speed of the canon shell fired at B so that the canon shells collide each other?

OpenStudy (amistre64):

determine the time it takes for A to get there, then use that to adjust for B

OpenStudy (amistre64):

A = gt^2/2 + 100 sin(30) t, 100cos(30)t = 700 solve for t in the distance to define height and time to get to B

OpenStudy (amistre64):

and i forgot the negative in front for A but ... if we assume that we want the shot to hit on the upswing, and not the downswing ... B = -gt^2/2 + vt = height of A solve for v